OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=?

OpenStudy (anonymous):

I HAVE AN ANSWER

OpenStudy (anonymous):



OpenStudy (anonymous):

\[a^2+b^2=c^2\] \[6^2+2^2=c^2\] etc yay! i hope it involves a square root

OpenStudy (anonymous):

you got it !